# What Makes a Visualization Complex?

### v1

@@ Acknowledgements.
- [110] G. Ramanarayanan, K. Bala, J. A. Ferwerda, and B. Walter. Dimensionality of visual complexity in computer graphics scenes. In Hum Vis Electron Imaging XIII, vol. 6806, pp. 142–151, 2008. doi: 10/fg9chj
- [111] A. R. Rao and G. L. Lohse. Towards a texture naming system: Identifying relevant dimensions of texture. Vision Research, 36(11):1649–1669, 1996. doi: 10/fn7t6w
- [112] K. Reinecke, T. Yeh, L. Miratrix, R. Mardiko, Y. Zhao, J. Liu, and K. Z. Gajos. Predicting users’ first impressions of website aesthetics with a quantification of perceived visual complexity and colorfulness. In Proc. SIGCHI Conf., pp. 2049–2058, 2013. doi: 10/gfs76j
- [113] R. A. Rensink. Visualization as a stimulus domain for vision science. J Vision, 21(8):3–3, 2021. doi: 10/mg29
- [114] I. Reppa, D. Playfoot, and S. J. McDougall. Visual Aesthetic Appeal Speeds Processing of Complex but not Simple Icons. In Proc. Hum. Factors Ergon. Soc. Annu. Meet., vol. 52(18), pp. 1155–1159, 2008. doi: 10/fx6f5r
- [115] J. Rigau, M. Feixas, and M. Sbert. Conceptualizing birkhoff’s aesthetic measure using shannon entropy and kolmogorov complexity. In Comput. Aesthet. Graph. Vis. Imaging, pp. 105–112, 2007. doi: 10/m9kx
@@
- [119] R. Rosenholtz, Y. Li, and L. Nakano. Feature congestion and subband entropy measures of visual clutter. J Vision, 6(6):827–827, 2006.
@@
- [124] E. Saraee, M. Jalal, and M. Betke. Visual complexity analysis using deep intermediate-layer features. Computer Vision and Image Understanding, 195:102949, 2020. doi: 10/gpzpdb
- [125] K. Sarıtaş, P. Dayan, K. Shen, and S. S. Nath. Complexity in complexity: Understanding visual complexity through structure, color, and surprise. In 2nd on Representational Alignment at ICLR, 2025.
@@
- [128] S. Silva, B. S. Santos, and J. Madeira. Using color in visualization: A survey. Comput Graph, 35(2):320–333, 2011. doi: 10/b59k9m
@@
- [130] S. Sonkusare, M. Breakspear, and C. Guo. Naturalistic stimuli in neuroscience: critically acclaimed. Trends in cognitive sciences, 23(8):699–714, 2019. doi: 10/gf4r7b
@@
- [134] A. N. Tuch, J. A. Bargas-Avila, K. Opwis, and F. H. Wilhelm. Visual complexity of websites: Effects on users’ experience, physiology, performance, and memory. IJHCS, 67(9):703–715, 2009. doi: 10/fk2j8c
- [135] E. R. Tufte. The visual display of quantitative information, vol. 2. Graphics press Cheshire, CT, 2001. doi: 10/bm26pp
@@
- [139] S. Wallace, R. Treitman, N. Kumawat, K. Arpin, J. Huang, B. Sawyer, and Z. Bylinskii. Towards readability individuation: the right changes to text format make large impacts on reading speed. J Vision, 20(10):17–17, 2020. doi: 10.1145/1235
@@
- [143] M. Wattenberg. A note on space-filling visualizations and space-filling curves. In IEEE INFOVIS, pp. 181–186, 2005. doi: 10/ghgt7p
@@
- [150] F. Windhager, A. Abdul-Rahman, M.-J. Bludau, N. Hengesbach, H. Lamqaddam, I. Meirelles, B. Speckmann, and M. Correll. Complexity as design material position paper. In IEEE BELIV, pp. 71–80, 2024. doi: 10/nvwm
- [151] J. M. Wolfe. Guided Search 6.0: An updated model of visual search. Psychon. Bull. Rev., 28(4):1060–1092, 2021. doi: 10/gh2s45
- [152] L. H. Wurm, G. E. Legge, L. M. Isenberg, and A. Luebker. Color improves object recognition in normal and low vision. J. Exp. Psychol. Hum. Percept. Perform., 19(4):899, 1993.
@@
- [156] V. Yoghourdjian, D. Archambault, S. Diehl, T. Dwyer, K. Klein, H. C. Purchase, and H.-Y. Wu. Exploring the limits of complexity: A survey of empirical studies on graph visualisation. Vis Inf, 2(4):264–282, 2018. doi: 10/ggjnn5
@@
- [159] Y. Zhao, H. Jiang, Q. Chen, Y. Qin, H. Xie, Y. Wu, S. Liu, Z. Zhou, J. Xia, and F. Zhou. Preserving minority structures in graph sampling. IEEE TVCG, 27:1698–1708, 2020. doi: 10/gt6nb8
@@
What Makes a Visualization Complex?
@@
Our perceived visual complexity (VC) dataset, VisComplexity2K, consists of 1,800 visualization images, each annotated with perceived VC scores collected from human participants via the online crowdsourcing platform Prolific. The dataset also includes corresponding objective image-quality metrics and associated metadata. VisComplexity2K is publicly available through a Google spreadsheet go.osu.edu/vcdatalink. All data can be accessed on tinyurl.com/vcgoogledrive, as well as on osf.io/bdet6. The key columns of VisComplexity2K include:
@@
The thumbnail of each image (column U), either the original image or a sub-panel from the source image, provides a gateway for image annotation and analysis, e. g., through Google CoLab run-time execution go.osu.edu/vccolablink.
@@
An indicator (column J) specifying the HP-Trees and Networks visualization images, which are the direct node-link diagrams used in Purchase et al. [107] and the RR-Color and Textured images, representing the texture and color stimuli used in Rosenholtz et al. [120] to support the analysis in the main text.
@@
The image link (column K) that points to a web storage address where a full-resolution version is accessible through tinyurl.com/viscomplexityimage. Please note that all VIS30K image files are copyrighted, and for most the copyright is owned by IEEE. Some have creative commons licenses or are in the public domain. Yet other images are subject to different, specific copyrights as indicated in the figure caption in the paper.
@@
The perceived visual complexity (column M) for the images collected through the crowdsourcing. The interface to collect this data is shown in Apdx. Figure 9.
@@
The O.MeC score (column BN) and The O.TiR score (column BJ), corresponding to the number of color and text-ink-ratio, respectively. The Python code for computing these is available at tinyurl.com/MeCTiRcomputation.
@@
The other 10 image-metrics, as shown in column M to BE. The Python code for computing these is available at tinyurl.com/metricscomputation.
@@ Appendix B Figure Image Courtesy by References
Since IEEE VIS 2025 has 2-page citation limits and we used many images from publications, we provide this figure-by-figure citation list to credit the image source. Images in What Makes a Visualization Complex? are taken from [11, 28, 123, 42, 64, 154, 56, 102, 69, 54, 121, 137, 78, 84, 39, 23, 149, 71]. Images in Figure 3 are taken from [11, 12]. Images in Figure 4 are taken from [146, 108, 11]. Images in Figure 7 are taken from [108, 81]. Images in Figure 8 are taken from [64, 88, 157, 143].
@@ C.1 Experimental Data Samples
The main criteria when we chose the data was diversity of Stimuli. Large, heterogeneous dataset spanning multiple visualization (node-link diagrams, heatmaps, charts), drawn from real-world sources, in order to increase ecological validity and generalizability of findings.
@@ C.2 A Workshop and Pilot Studies
The second approach involved a hierarchical division of images into groups. For example, in the study by Oliva et al. [99], participants divided 100 images into four hierarchies and stated their criteria to categorize natural scenes as ‘openness’ etc. This is also the standard de facto approach to measure perceptual dimensions from images (e. g., seminar work of texture and visualization image classifications [111, 83]). We piloted this approach in our workshop and found it effective for gathering participants’ comments to guide subsequent steps. However, similar to the first approach, it presented scalability challenges, limiting its applicability for larger datasets. During the workshop experiment, 46 participants, organized into 20 groups of 2-3 students, were instructed to perform the hierarchical sorting tasks based on visual complexity, defined as “the amount of detail or intricacy in images” [107]. These participants were graduate students having domain expertise in machine learning and computer vision, before they took the visualization class; and three participants had extensive industrial experience. Participants were asked to document the factors that influenced their sorting decisions. Participants on day 1 sorted 20 images, with all groups completing the task in an average of 20 minutes.
The scalability challenge arrived on day 2 for sorting 50 images, with the same instruction. Participants described the sorting task as "overwhelming." No group was able to finish within 1.5 hours, and only of the groups managed to complete the task within 2 hours. First, participants noted that increasing image set made any linear sorting impractical; and two groups utilized a 2D Cartesian coordinate system to organize VC clusters. Second, their comments suggested the benefits of text display: without the knowledge of the data source and differences between MASSVIS and VIS30K, 11 groups pointed out that some (MASSVIS) images ‘looked nice’ and these (VIS30K) show sophisticated scientific phenomena. ‘Text (annotated) greatly improved clarity; thus I rated the image with lower complexity’. Participants also commented that ‘After reading the texts, I know where to look.’ Others mentioned object set-size and said ‘there were not many items in the image; thus these (MASSVIS) images were simpler.’ Third, half of the groups commented that “these (VIS30K scientific visualization) images were difficult to understand without text or legend showing what an image represents.” These subjective experiences led us to exclude 3D surface and volume rendering visualizations from our formal experiment, as they were challenging for participants and were likely to be rated more complex.
@@
We finally used a multi-stage active sampling algorithm [90], which reduces the number of comparisons by 10 folds while maintaining the same level of reliability. In each phase, the algorithm optimizes the information obtained to adapt the algorithms to assign an absolute score. Figure 10 shows the validation of this approach.
@@ C.3 Partial Least Square (PLS) for Variable Selection: Additional Discussion
PLS is a relatively new method and is gaining interests in choosing variables in many scientific domains such as bio-informatics, chemistry and neuroscience [73],mainly due to scientists’ ability to collect large datasets with many variables and PLS’s ability to handle a large number of variables. The process of choosing the most useful variable is called variable selection, where variables in our case are the objective image metrics (predictors) and the response is the collected VC scores via the crowdsourcing experiment. Unlike principal component regression, which maximizes variance in predictors, PLS explicitly maximizes the covariance between predictors and the target variable (VC score), allowing us to extract latent components that are optimally aligned with perceptual responses. This made PLS an appropriate tool to retain and analyze all metrics jointly while still yielding interpretable component structures. We further incorporated a bootstrapping procedure to assess the reliability of each metric’s contribution. Specifically, we resampled the data with replacement across multiple iterations to estimate the distribution of the PLS regression coefficients, from which we computed standard errors and approximate p-values. This enabled us to identify statistically significant and stable metric variables.
Our calculation. We assessed the appropriate number of PLS components to retain by running models with 1 to 10 components and selecting the number where the regression began to plateau. Based on this procedure across different experiments, we selected 5 components for all final analyses to ensure a balance between model complexity and interpretability. Lastly, to aid interpretation, we complemented coefficient analysis with individual metric vs. VC plots (e. g., Figure 6 and extended in the supplemental) that visualize how VC scores vary across the value range of each feature.
@@ C.4 Verbal Reasoning
While the main text focuses on quantitative analysis, we also conducted an additional analysis of the 698 self-reported verbal reasoning responses collected from the post-questionnaire. In these responses, participants explained their rationale for two randomly selected trials. The words and their frequencies offer insights into the cognitive processes involved, helping to construct vocabularies that represent semantic categories relevant to the perception of visual complexity. This approach aligns with prior work in high-order perception research, such as studies on memorability [72] and attitude [4]. Table 6 shows the proportion of words grouped by semantic category, ordered by descending word frequency. Table 7 summarizes the resulting comments into the categories. The most frequently cited themes were clarity and readability () and color and contrast (), followed by information density (), number of elements (), abstractness and familiarity (), visual clutter (), interconnectedness (), and beautifulness ().
@@
The high-level attributes of clarity and readability associated with verbal reasoning may explain that some high-level details could not be explained by the PLS models. To communication purposes, the familiarity, structure, surprise [125], distinctiveness [74, 75], aesthetic appeal, are important. To domain scientists which we did not measure, correctness (i. e., the representations to use visualization that represent key knowledge in order to access meaning) and semantic distance (the goodness-of-fit between representation and its meaning), also matter. The image-based metrics do not represent this set of high-level attributes but these can also be captured using our method.

### v3

@@ Acknowledgements.
- [104] R. Y. Pang, K. Feng, S. Feng, C. Li, W. Shi, Y. Tsvetkov, J. Heer, and K. Reinecke. Interactive reasoning: Visualizing and controlling chain-of-thought reasoning in large language models. arXiv preprint arXiv:2506.23678, 2025. doi: 10/pzxr
- [105] J. Perkiö and A. Hyvärinen. Modelling image complexity by independent component analysis, with application to content-based image retrieval. In International Conference on Artificial Neural Networks, pp. 704–714. Springer, 2009. doi: 10/cm97t9
@@
- [110] G. Ramanarayanan, K. Bala, J. A. Ferwerda, and B. Walter. Dimensionality of visual complexity in computer graphics scenes. In Hum Vis Elec Imaging XIII, vol. 6806, pp. 142–151, 2008. doi: 10/fg9chj
- [111] A. R. Rao and G. L. Lohse. Towards a Texture Naming System: Identifying relevant dimensions of texture. Vision Res, 36(11):1649–1669, 1996. doi: 10/fn7t6w
- [112] K. Reinecke, T. Yeh, L. Miratrix, R. Mardiko, Y. Zhao, J. Liu, and K. Z. Gajos. Predicting users’ first impressions of website aesthetics with a quantification of perceived visual complexity and colorfulness. In ACM SIGCHI, pp. 2049–2058, 2013. doi: 10/gfs76j
- [113] R. A. Rensink. Visualization as a stimulus domain for vision science. J Vision, 21(3):1–18, 2021. doi: 10/mg29
- [114] I. Reppa, D. Playfoot, and S. J. McDougall. Visual aesthetic appeal speeds processing of complex but not simple icons. In Proc Hum Factors Ergon Soc Annu Meet, vol. 52(18), pp. 1155–1159, 2008. doi: 10/fx6f5r
- [115] J. Rigau, M. Feixas, and M. Sbert. Conceptualizing birkhoff’s aesthetic measure using shannon entropy and kolmogorov complexity. In Comput Aesthet Graph Vis Imaging, pp. 105–112, 2007. doi: 10/m9kx
@@
- [117] B. E. Rogowitz, T. Frese, J. R. Smith, C. A. Bouman, and E. B. Kalin. Perceptual image similarity experiments. In Hum Vis Elec Imaging VIII, vol. 3299, pp. 576–590. SPIE, 1998. doi: 10/cn92mg
@@
- [124] E. Saraee, M. Jalal, and M. Betke. Visual complexity analysis using deep intermediate-layer features. Comput Vis Img Understanding, 195:102949, 2020. doi: 10/gpzpdb
- [125] K. Sarıtaş, P. Dayan, K. Shen, and S. S. Nath. Complexity in Complexity: Understanding visual complexity through structure, color, and surprise. In 2nd on Representational Alignment at ICLR, 2025. doi: 10/pzj3
@@
- [127] C. E. Shannon. A mathematical theory of communication. The Bell system technical journal, 27(3):379–423, 1948.
@@
- [129] S. Silva, B. S. Santos, and J. Madeira. Using color in visualization: A survey. Comput & Graph, 35(2):320–333, 2011. doi: 10/b59k9m
@@
- [131] S. Sonkusare, M. Breakspear, and C. Guo. Naturalistic Stimuli in Neuroscience: Critically acclaimed. Trends in Cog Sci, 23(8):699–714, 2019. doi: 10/gf4r7b
@@
- [133] M. Stone and L. Bartram. Alpha, contrast and the perception of visual metadata. In Color and Imaging Conf, vol. 16, pp. 355–359, 2008. doi: 10/pzvz
@@
- [136] D.-Y. Tsai, Y. Lee, and E. Matsuyama. Information entropy measure for evaluation of image quality. Journal of digital imaging, 21(3):338–347, 2008. doi: 10/fxgzpx
- [137] A. N. Tuch, J. A. Bargas-Avila, K. Opwis, and F. H. Wilhelm. Visual complexity of websites: Effects on users’ experience, physiology, performance, and memory. Int J Hum Comput Stud, 67(9):703–715, 2009. doi: 10/fk2j8c
@@
- [141] S. Wallace, Z. Bylinskii, J. Dobres, B. Kerr, S. Berlow, R. Treitman, N. Kumawat, K. Arpin, D. B. Miller, J. Huang, et al. Towards individuated reading experiences: Different fonts increase reading speed for different individuals. ACM Trans Comput Hum Interact, 29(4):1–56, 2022. doi: 10/gqg7jp
@@
- [145] M. Wattenberg. A note on space-filling visualizations and space-filling curves. In IEEE InfoVis, pp. 181–186, 2005. doi: 10/dsngsw
@@
- [152] F. Windhager, A. Abdul-Rahman, M.-J. Bludau, N. Hengesbach, H. Lamqaddam, I. Meirelles, B. Speckmann, and M. Correll. Complexity as design material position paper. In IEEE BELIV Workshop, pp. 71–80, 2024. doi: 10/nvwm
- [153] J. M. Wolfe. Guided Search 6.0: An updated model of visual search. Psychol Bull Rev, 28(4):1060–1092, 2021. doi: 10/gh2s45
- [154] L. H. Wurm, G. E. Legge, L. M. Isenberg, and A. Luebker. Color improves object recognition in normal and low vision. J Exp Psychol Hum Percept Perform, 19(4):899, 1993. doi: 10/b9w9k9
@@
- [158] V. Yoghourdjian, D. Archambault, S. Diehl, T. Dwyer, K. Klein, H. C. Purchase, and H.-Y. Wu. Exploring the limits of complexity: A survey of empirical studies on graph visualisation. Visual Informatics, 2(4):264–282, 2018. doi: 10/ggjnn5
@@
- [161] Y. Zhao, H. Jiang, Q. Chen, Y. Qin, H. Xie, Y. Wu, S. Liu, Z. Zhou, J. Xia, and F. Zhou. Preserving minority structures in graph sampling. IEEE TVCG, 27(2):1698–1708, 2021. doi: 10/gt6nb8
@@
What Makes a Visualization Image Complex?
@@
Our perceived visual complexity (VC) dataset, VisComplexity2K, consists of 1,800 visualization images, each annotated with VC scores collected from human participants via the online crowdsourcing platform Prolific. The dataset also includes corresponding objective image-quality metrics and associated metadata. VisComplexity2K is publicly available through a Google spreadsheet go.osu.edu/viscomplexitydata. All data analysis code and images can be accessed via Colab on go.osu.edu/vcgoogledrive. osf.io/bdet6 also has all the analysis results and code. The key columns of VisComplexity2K include:
@@
The thumbnail of each image (column E), either the original image or a sub-panel from the source image, provides a gateway for image annotation and analysis, e. g., through Google CoLab run-time execution go.osu.edu/vcmodel.
@@
An indicator (column P) specifying the HP-Trees and Networks visualization images, which are the direct node-link diagrams used in Purchase et al. [107] and the RR-Color and Textured images, representing the texture and color stimuli used in Rosenholtz et al. [120] to support the analysis in the main text.
@@
The image link (column D) that points to a web storage address where a full-resolution version is accessible through go.osu.edu/vcgoogledrive. Please note that all VIS30K image files are copyrighted, and for most the copyright is owned by IEEE. Some have creative commons licenses or are in the public domain. Yet other images are subject to different, specific copyrights as indicated in the figure caption in the paper.
@@
The perceived visual complexity (column F) for the images collected through the crowdsourcing. The interface to collect this data is shown in Apdx. Figure 9.
@@
The O.MeC score (column BO) and The O.TiR score (column BJ), corresponding to the number of color and text-ink-ratio, respectively. The Python code for computing these is available in the sub-folders Metric.MeaningfulColor(O.MeC) and Metric.TexttoInkRatio(O.TiR) on go.osu.edu/vcgoogledrive.
@@
The other 10 image-metrics, as shown in column V to BF. The Python code for computing these is available in the sub-folder Metrics.10 on go.osu.edu/vcgoogledrive.
@@ Appendix B Figure Image Courtesy by References
Since IEEE VIS 2025 has a 2-page citation limits and we used many images from publications, we provide this figure-by-figure citation list to credit the image source. Images in What Makes a Visualization Image Complex? are taken from [10, 28, 123, 40, 62, 156, 54, 101, 67, 52, 121, 139, 76, 83, 38, 23, 151, 69]. Images in Figure 3 are taken from [10, 11]. Images in Figure 4 are taken from [148, 108, 10]. Images in Figure 7 are taken from [108, 80]. Images in Figure 8 are taken from [62, 87, 159, 145].
@@ C.1 Experimental Data Samples
The main criteria when we chose the data was diversity of stimuli. It is a large, heterogeneous dataset spanning multiple visualization (node-link diagrams, heatmaps, charts), drawn from real-world sources, in order to increase ecological validity and generalizability of findings.
@@ C.2 A Workshop and Pilot Studies
The second approach involved a hierarchical division of images into groups. For example, in the study by Oliva et al. [98], participants divided 100 images into four hierarchies and stated their criteria to categorize natural scenes as ‘openness’ etc. This is also the standard de facto approach to measure other perceptual dimensions from images (e. g., seminal work of texture and visualization image classifications [111, 82]). We piloted this approach in our workshop and found it effective for gathering participants’ comments to guide subsequent steps. However, similar to the first approach, it presented scalability challenges, limiting its applicability for larger datasets. During the workshop experiment, 46 participants, organized into 20 groups of 2-3 students, were instructed to perform the hierarchical sorting tasks based on perceived visual complexity of the images, defined as “the amount of detail or intricacy in images” [107]. These participants were graduate students having domain expertise in machine learning and computer vision, before they took the visualization class; and three participants had extensive industrial experience. Participants were asked to document the factors that influenced their sorting decisions.
Participants on day 1 sorted 20 images, with all groups completing the task in an average of 20 minutes. The scalability challenge arrived on day 2 for sorting 50 images, with the same instruction. Participants described the sorting task as "overwhelming." No group was able to finish within 1.5 hours, and only of the groups managed to complete the task within 2 hours. First, participants noted that increasing image set made any linear sorting impractical; and two groups utilized a 2D Cartesian coordinate system to organize VC clusters. Second, their comments suggested the benefits of text display: without the knowledge of the data source and differences between MASSVIS and VIS30K, 11 groups pointed out that some (MASSVIS) images ‘looked nice’ and these (VIS30K) show sophisticated scientific phenomena. ‘Text (annotated) greatly improved clarity; thus I rated the image with lower complexity’. Participants also commented that ‘After reading the texts, I know where to look.’ Others mentioned object set-size and said ‘there were not many items in the image; thus these (MASSVIS) images were simpler.’ Third, half of the groups commented that “these (VIS30K scientific visualization) images were difficult to understand without text or legend showing what an image represents.” These subjective experiences led us to exclude 3D surface and volume rendering visualizations from our formal experiment, as they were challenging for participants and were likely to be rated more complex.
@@
We finally used a multi-stage active sampling algorithm [89], which reduces the number of comparisons by 10 fold while maintaining the same level of reliability. In each phase, the algorithm optimizes the information obtained to adapt the algorithms to assign an absolute score. Figure 10 shows the validation of this approach.
@@ C.3 Partial Least Square (PLS) for Variable Selection: Additional Discussion
PLS is a relatively new method and is gaining interests in choosing variables in many scientific domains such as bio-informatics, chemistry and neuroscience [71]. It is becoming widely used mainly due to scientists’ ability to collect large datasets with many variables and PLS’s ability to handle a large number of variables. The process of choosing the most useful variable is called variable selection, where variables in our case are the objective image metrics (predictors) and the response is the collected VC scores via the crowdsourcing experiment. Unlike principal component regression, which maximizes variance in predictors, PLS explicitly maximizes the covariance between predictors and the target variable (VC score), allowing us to extract latent components that are optimally aligned with perceptual responses. This made PLS an appropriate tool to retain and analyze all metrics jointly while still yielding interpretable component structures.
Our calculation. We assessed the appropriate number of PLS components to retain by running models with 1 to 10 components and selecting the number where the regression began to plateau. Based on this procedure across different experiments, we selected 5 components for all final analyses to ensure a balance between model complexity and interpretability. For the bootstrapping procedure to assess the reliability of each metric’s contribution, we resampled the data with replacement across multiple iterations to estimate the distribution of the PLS regression coefficients, from which we computed standard errors and approximate p-values. This enabled us to identify statistically significant and stable metric variables. Lastly, to aid interpretation, we complemented coefficient analysis with individual metric vs. VC plots (e. g., Figure 6 and extended in the supplemental) that visualize how VC scores vary across the value range of each feature.
@@ C.4 Verbal Reasoning
While the main text focuses on quantitative analysis, we also conducted an additional analysis of the 698 self-reported verbal reasoning responses collected from the post-questionnaire. In these responses, participants explained their rationale for two randomly selected trials. The words and their frequencies offer insights into the cognitive processes involved, helping to construct vocabularies that represent semantic categories relevant to the perception of visual complexity.
This approach aligns with prior work in high-order perception research, such as studies on memorability [70] and attitude [4]. Table 6 shows the proportion of words grouped by semantic category, ordered by descending word frequency. Table 7 summarizes the resulting comments into the categories. The most frequently cited themes were clarity and readability () and color and contrast (), followed by information density (), number of elements (), abstractness and familiarity (), visual clutter (), interconnectedness (), and beautifulness ().
@@
The high-level attributes of clarity and readability associated with verbal reasoning may explain that some high-level details could not be explained by the PLS models. For communication purposes, the familiarity, structure, surprise [125], distinctiveness [72, 73], and aesthetic appeal, are important. Furthermore, we also did not measure correctness, (i. e., the representations to use visualization that represent key knowledge in order to access meaning) and semantic distance (the goodness-of-fit between representation and its meaning). Our image-based metrics do not represent this set of high-level attributes, but our method can capture these to further examine complexity.